You decide to start your own company and borrow $750000. You plan to fully amortize one-half of the loan with a balloon payment at the end of the term of the loan in 5 years. What would be the annual payment? When the loan matures, how much must you pay the lender if the interest rate on the loan is 5.75%(what is final payment)?

Calculating with BA II Financial Calculator
Answer #1

Loan is paid in two parts

a) Half loan amount is amortised i.e. $375000

b) Balance half loan amount is paid at the end of 5th year

Calculation of Annual Payment

Annual Installment for one-half loan= {Loan Amount/Pvaf(i,n)}

[375000 / {1/(1.0575)^5)}]

$88419.05

Therefore annual payment= $88419.05

Calculation for Final Payment on maturity of loan

i) For Balance half loan amount = $375000

Interest=5.75%

Amount after 5 years= [375000*{(1.0575)^5}] = $495944.58

ii) Fifth Installment amount as calculated above = $88419.05

Hence Final payment = i + ii = $584363.63